Bonus homework for class Please show Formulas! thanks in advance!

1. Calculating OCF [LO1] Consider the following income statement:
Sales $713,500
Cost 497,300
EBIT ?
Taxes (34%) ?
Net Income ?
Fill in the missing numbers and then calculate the OCF. What is the depreciation tax shield?

2. Operating Cash Flow and Leverage [LO41 A proposed project has fixed costs of $83,000 per year. The operating cash flow at 9,100 units is $102,900. Ignoring the effect of taxes, what is the degree of operating leverage? If units sold rise from 9,100to 9,500, what will be the increase in operating cash flow? What is the new degree of operating leverage?
Answer #1

1) EBIT = Sales - Cost = 713,500 - 47,300 = $216,200

Taxes = Earnings before tax*tax rate = 216,200 * 34% = $73,508

Net Income = Earnings before tax - tax amount = 216,200 - 73,508 = $142,692

OCF or Operating Cash Flow = Net income + depreciation and other non-cash expenses - net changes in working capital

Since the information about depreciation and net changes in working capital is not given,

OCF = Net Income = $142,692

For tax purposes, depreciation is deducted from taxable income, thus reducing the income on which tax is calculated. This reduces the tax payable amount. So, the depreciation tax shield is the amount by which the depreciation shields or protects the taxpayer from paying income taxes, i.e., Depreciation tax shield = Depreciation * Tax Rate

2)

a) Fixed Cost = $83,000

Operating cash flow at 9100 units = $102,900

Total cost = Fixed Cost + Variable Cost

Profit (operating cash flow) = Sales - Total Cost = Sales - (Fixed Cost + Variable Cost)  

So, 102,900 = Sales - (83,000 + Variable Cost) = Sales - 83,000 - Variable Cost

Sales - Variable Cost = $185,900

Contribution margin = Sales - Variable Cost

So, Contribution Margin = $185,900

Contribution Margin per unit = Contribution Margin / Number of units = 185,900 / 9100 = $20.4286

Degree of Operating Leverage = Contribution margin / Profit = 185,900 / 102,900 = 1.8066

b) Units increase from 9100 to 9,500

Contribution Margin Per unit = $20.4286

Contribution Margin = Contribution Margin Per unit * Number of units = 20.4286 * 9500 = $194,072

Operating Cash flow = Sales - Fixed Cost - Variable Cost

Operating Cash flow = (Sales - Variable Cost) - Fixed Cost

Operating Cash flow = Contribution margin - Fixed Cost = 194,072 - 83,000 = $111,072

So, if units sold increase from 9100 to 9500, operating cash flow increases from $102,900 to $111,072 (by $8,172).

New Degree of Operating Leverage = Contribution margin / Profit = 194,072 / 111,072 = 1.7473